question: A box contains 3 red, 7 white & 2 blue marbles. suppose 3 marbles are chosen with replacement. Determine the probability that the selected marbles are different colors.
Answer #1

Given that, Box contains 3 red, 7 white & 2 blue marbles. Total no. of marbles in a box is 3+7+2= 12 marbles

We have to find the probability that the selected marbles are different colours i.e. there should be 1 red, 1 white & 1 blue.

required prob.= (3C1 * 7C1 * 2C1)/ 12C3= (3*7*2)/220= 42/220= 0.191